OpenStudy (anonymous):

Multiply two and four-fifths negative five and two-thirds . negative ten and eight-fifteenths negative fifteen and thirteen-fifteenths ten and eight-fifteenths fifteen and thirteen-fifteenths

OpenStudy (alexandervonhumboldt2):

2*4/5=8/5

OpenStudy (alexandervonhumboldt2):

-5*2/3=-10/3

OpenStudy (alexandervonhumboldt2):

-(3*4*2)=-24

OpenStudy (anonymous):

@AlexandervonHumboldt2 i don't know which one